# Coloring minimal Cayley graphs

### v2

@@ 1. Introduction
Babai conjectured in 1994 a negative answer even in the case the group is finite. Note that assuming minimality here is essential. Indeed, every graph is an induced subgraph of some Cayley graph [6, 17, 4], this is far from being the case for minimal Cayley graphs, which are known to be sparse [3, 32]. However, there also exist (non-minimal) sparse Cayley graphs (of arbitrary girth) that are expanders and hence have unbounded chromatic number [23]. The chromatic number of random Cayley graphs has been studied by Alon [2] and Green [18] as well as some concrete Cayley graphs have been considered recently [9, 10]. However, none of these are minimal. Concerning the chromatic number of minimal Cayley graphs, in [5, Section 3.4] Babai mentions the following:
@@ 2. Positive results
Let us begin with the question whether every finite group admits a minimal Cayley graph of bounded chromatic number. This was already answered by Babai [3, 5]. Indeed, he proved that every finite group has a 3-colorable minimal Cayley graph. His proof depends on the fact that every finite simple group can be generated by two elements, one of which is an involution. This is a known consequence of the Classification of Finite Simple Groups. For completeness we include a proof following Babai’s lines. At the basis of this we have the following [3, Lemma 4.2], which will also use later on. For a group $\Gamma$ denote its minimum chromatic number and $\chi_{\mathrm{min}}(\Gamma)$ the minimum $\chi(\mathrm{Cay}(\Gamma,C))$ over all generating sets $C$ of $\Gamma$.
@@
We can however use the above lemma to bound the maximum chromatic in some cases. A group is called Dedekind if all its subgroups are normal. Clearly, this includes all abelian groups, and by results of Dedekind [11] and Baer [7], actually not much more.
@@
Let $G={\rm Cay}(\Gamma,C)$ be the Cayley graph of $\Gamma$ with respect to a minimal set of generators $C$. Since $\Gamma^{\prime}\subseteq\Phi(\Gamma)$, then it follows that $\Gamma/\Phi(\Gamma)$ is commutative and minimally generated by $C/\Phi(\Gamma)$. Then, by Theorem 2.5, ${\rm Cay}(\Gamma/\Phi(\Gamma),C/\Phi(\Gamma))$ is $3$-colorable. Since $C\subset(C/\Phi(\Gamma))\cdot\Phi(\Gamma)$, by Lemma 2.1 we conclude that ${\rm Cay}(\Gamma,C)$ is $3$-colorable. ∎
@@
$W_{b}(n)=\log_{2}(\frac{n}{W_{b}(n)})=\log_{2}(\frac{n}{\log_{2}(\frac{n}{W_{b}(n)})}).$
@@
$W_{b}(n)<\log_{2}(\frac{n}{\log_{2}(\frac{n}{\log_{2}(n)})})=\log_{2}n-\log_{2}\log_{2}(\frac{n}{\log_{2}n}),$
@@ 3. Lower bounds
We begin this section by presenting a minimal Cayley graph of chromatic number $4$ and a semiminimal Cayley graphs of chromatic number $7$.
@@
We proceed by induction on $k\geq 1$. Chose $G_{1}$ to be the graph with a single vertex and $G_{2}=C_{4}$ with edges colored alternatingly with two colors. If $k\geq 3$ then denote by $n$ the number of vertices of $G_{k-1}$. Define $X$ as a set of $(k-1)(n-1)+1$ new vertices without edges. Now, for every subset of $Y$ size $n$ of $X$ take a copy $G^{\prime}$ of $G_{k-1}$ (where all copies can be considered to be edge-colored with the same set) and add a matching of size $n$ that connects the vertices of $Y$ with the vertices of $G^{\prime}$. Each of these new matchings will be edge colored with its own private color. The resulting graph is $G_{k}$.
@@
While we have negatively answered the one-popular-color variant of Question 1.3 its no-lonely-color variant remains open. Let us propose a strengthening of it.
@@
Note that this conjecture for $d=2$ implies a positive answer to the weak variant of Question 1.3. However, it is open even for the case $d=1$. On the other hand, it also remains open if graphs with a one-popular-color coloring exist that satisfy Conjecture . Finally, note that the questions about the boundedness of the chromatic number of minimal or semiminimal Cayley graphs of general finite groups remain open.
@@ 4. Acknowledgements
We thank Laszlo Babai for several comments on an earlier version of this paper.

### v3

@@ 1. Introduction
Babai conjectured in 1994 a negative answer even in the case the group is finite. Note that assuming minimality here is essential. Indeed, every graph is an induced subgraph of some Cayley graph [6, 17, 4], this is far from being the case for minimal Cayley graphs, which are known to be sparse [3, 32]. However, there also exist (non-minimal) sparse Cayley graphs (of arbitrary girth) that are expanders and hence have unbounded chromatic number [23]. The chromatic number of random Cayley graphs has been studied by Alon [2] and Green [18] and some concrete Cayley graphs have been considered recently [9, 10]. However, none of these are minimal. Concerning the chromatic number of minimal Cayley graphs, in [5, Section 3.4] Babai mentions the following:
@@ 2. Positive results
Let us begin with the question whether every finite group admits a minimal Cayley graph of bounded chromatic number. This was already answered by Babai [3, 5]. Indeed, he proved that every finite group has a 3-colorable minimal Cayley graph. His proof depends on the fact that every finite simple group can be generated by two elements, one of which is an involution. This is a known consequence of the Classification of Finite Simple Groups. For completeness we include a proof following Babai’s lines. At the basis of this we have the following [3, Lemma 4.2], which we will also use later on. For a group $\Gamma$ denote its minimum chromatic number and $\chi_{\mathrm{min}}(\Gamma)$ the minimum $\chi(\mathrm{Cay}(\Gamma,C))$ over all generating sets $C$ of $\Gamma$.
@@
We can however use the above lemma to bound the maximum chromatic number in some cases. A group is called Dedekind if all its subgroups are normal. Clearly, this includes all abelian groups, and by results of Dedekind [11] and Baer [7], actually not much more.
@@
Let $G={\rm Cay}(\Gamma,C)$ be the Cayley graph of $\Gamma$ with respect to a minimal set of generators $C$. Since $\Gamma^{\prime}\subseteq\Phi(\Gamma)$, then it follows that $\Gamma/\Phi(\Gamma)$ is commutative and minimally generated by $C/\Phi(\Gamma)$. Then, by Theorem 2.5, ${\rm Cay}(\Gamma/\Phi(\Gamma),C/\Phi(\Gamma))$ is $3$-colorable. But now we are done because since $C\subset(C/\Phi(\Gamma))\cdot\Phi(\Gamma)$, by Lemma 2.1 we have that ${\rm Cay}(\Gamma,C)$ is $3$-colorable. ∎
@@
$W_{b}(n)=\log_{2}\left(\frac{n}{W_{b}(n)}\right)=\log_{2}\left(\frac{n}{\log_{2}\frac{n}{W_{b}(n)}}\right).$
@@
$W_{b}(n)<\log_{2}\left(\frac{n}{\log_{2}\frac{n}{\log_{2}(n)}}\right)=\log_{2}n-\log_{2}\log_{2}\left(\frac{n}{\log_{2}n}\right),$
@@ 3. Lower bounds
We begin this section by presenting a minimal Cayley graph of chromatic number $4$ and a semiminimal Cayley graph of chromatic number $7$.
@@
We proceed by induction on $k\geq 1$. Choose $G_{1}$ to be the graph with a single vertex and $G_{2}=C_{4}$ with edges colored alternatingly with two colors. If $k\geq 3$ then denote by $n$ the number of vertices of $G_{k-1}$. Define $X$ as a set of $(k-1)(n-1)+1$ new vertices without edges. Now, for every subset $Y$ of size $n$ of $X$ take a copy $G^{\prime}$ of $G_{k-1}$ (where all copies can be considered to be edge-colored with the same set) and add a matching of size $n$ that connects the vertices of $Y$ with the vertices of $G^{\prime}$. Each of these new matchings will be edge colored with its own private color. The resulting graph is $G_{k}$. Note that the choice of the matchings in the construction is not unique.
@@
While we have negatively answered the one popular color variant of Question 1.3 its no lonely color variant remains open. Let us propose a strengthening of it.
@@
Note that this conjecture for $d=2$ implies a positive answer to the weak variant of Question 1.3. However, it is open even for the case $d=1$. On the other hand, it also remains open if graphs with a one popular color coloring exist that satisfy Conjecture 1.2. Finally, note that the questions about the boundedness of the chromatic number of minimal or semiminimal Cayley graphs of general finite groups remain open.
@@ 4. Acknowledgements
We thank László Babai for several comments on an earlier version of this paper.